OpenStudy (anonymous):

Let f(x) =the square root of x, g(x) = x + 8 and h(x) = 9x. Find the domain of the following function then simplify the expression. (g − h)(x)

OpenStudy (anonymous):

(g-h)(x) = x+8-9x=8-8x=8(1-x) the domain is all real numbers

OpenStudy (anonymous):

oh wait sorry h=9/x !!

OpenStudy (anonymous):

ok similarly (g-h)(x) = x+8 - 9/x and the domain R−{0} (all real numbers except for x=0)

OpenStudy (anonymous):

so itd be x-1/x ?

OpenStudy (anonymous):

no, x +8 -9/x cannot be simplifed any further

OpenStudy (anonymous):

oh ok thank you!